Find the value of the current i in milliamps. R1 is 421 ohms. R2 is 433.5 ohms, v is 8 volts. vx is 1.8v1.

Find the value of the current i in milliamps. R1 i
Answer #1

According to Kirchoff's law:

v - iR1 + vx - iR2 = 0

v + 1.8v1 = i(R1 + R2)

8 + 1.8v1 = 854.5i   \small \rightarrow eq.1

v - v1 + vx - iR2 = 0

v - v1 + 1.8v1 = iR2

8 + 0.8v1 = 433.5i \small \rightarrow eq.2

9*(eq.2) - 4*(eq.1) \small \Rightarrow 438.5i = 40

i = 0.0827 A = 82.7 mA